OpenStudy (anonymous):

Help! Simplify: (cubed root (x^12z^9)/(y^15))^4 A. x^11 z^-10 y^-16 B. x^7 z^(-27/4) y^(-45/4) C. x^8zy^-1 D. x^16 z^-12 y^-20 Thanks!

OpenStudy (rock_mit182):

dude :\[\sqrt[3]{\frac{ x ^{12} z^{9} }{( y^{15})^{4} }}\] is that the algebraic expression ?

OpenStudy (rock_mit182):

@kidfisto ?

OpenStudy (anonymous):

The ^4 is for the whole cubed root not just y^15 @rock_mit182

OpenStudy (anonymous):

\[(\sqrt[3]({x^{12}z^{-9}})/ {y^{15}})^4\] and the 9 is negative sorry about that

OpenStudy (rock_mit182):

that makes more sense

OpenStudy (rock_mit182):